Klapperich Lab:Notebook/Lab Meeting Notes/2008/10/07

From OpenWetWare
Jump to navigationJump to search
Owwnotebook icon.png Project name Report.pngMain project page
Resultset previous.pngPrevious entry      Next entryResultset next.png

7 October 2008 Lab Meeting Agenda

† Sonali's research report today.

† Announcements

  • A slow migration to 709 can now begin.
  • JZ and JD out next week at mTAS.

† Flu R01

  • More sample back? Can we put in 1.5 ml saline, collect in 3ml VTM tube directly? YES. Cathie will amend.
  • Plaque assay status. Redo . Too dilute.
  • Cathie STILL STILL has to submit the progress report to INSPIR.

† SEPSIS Project - CIMIT Bird Flu

  • Sample (raw pellets and film) from Ticona COC ordered.
  • B sub in blood is done. Lets talk about comepleting the gram positive plots so they "match."
  • SEM image analysis of pore size. Progress?
  • Run E. coli RT-PCR using primers for a genomic target.
  • Report DNA amounts along with the RT-PCR data for the Sepsis paper.
  • Tube meeting took place last week. More to come.

†RNA project

  • Jeff Braman will visit in early Dec.
  • SM and AC did the first experiments this week. RNA obtained. Looks clean.
  • Do dilutions. Send to Bioanalyzer.


  • Mark finish control circuit
  • Fluorescence microscopy available from 720. Try beads experiment. Update?
  • Cutter plotter arrived.
  • Jane, Cathie, Larry, Ranjith will meet soon to discuss scale up of sol gel in chip process.
  • Discuss dispensing from evap chip in a real world way.
  • Current design gives E.coli signal, clotting is still a problem
  • Change of design: funnel-shaped wells, smooth hole, make samples in pipette tips?

† Valve Array/Valve building (FJ)

  • Waiting for fixture still from Melissa. Hopefully this week.
  • Update: Monday 6 Oct Haven't heard anything from Shane yet. Attempting to contact again, so far without success.

† Fraunhofer: LOAC

  • Update?
  • Alex working on more toluene bonding issues.
  • Try Ticona material. (HM, JD, AG)
  • Still Still Still troubleshooting ??

† Biointerfaces group

  • Mold still not made. Mask potentially Wed AM. When mask is made.
  • Do preliminary E. coli experiment soon. Sonali will train.

† CIMIT- Colson Grant

  • Preliminary plan for IRB submission at Harvard in place. Target date 1/1/09
  • STAR-CD Renewal - January ($2K)


  • Physical PCR areas defined in 709 (check with Sonali).
  • QQ continued with Taqman assays.
  • QQ will run it on a gel to check.
  • Run Gels for water Wash steps - show me and Sonali
  • Only 1/5 chips are good at the hot embossing step. The vaccum chuck will be checked and o-rings changed - Will try after fixes and report next week.

* Fixuture for PCR chip. The heater and the thermocouples would be on the fixture.

  • MinCheol is making Four simple PCR models to optimize temperature distribution along to the channel. Will do another model with the insulation on the bottom.
  • MinCheol is simulating DNA's Residence Time for simple PCR models. The DNA particles need to be much smaller. Also, consult the Lab on Chip paper that JD mentioned.
  • Have a PCR only meeting sometime next week.
  • Qingqing has to do experiments with MinCheol's designs.


  • Cathie Still Still Working on Whitepaper almost done!

† Senior project

  • Megan ordering oligos and primers for the RCA. Quote. Check with Heiko on his sources.
  • Haven't seen Teddy yet. What is up with him?

† F31: Cochlea

  • ARO Abstract submitted.
  • Data summary focusing on the fundamental biological questions for next week.

† Silica Optimization (Lambda):

  • Ongoing. More concentrations. Forged chips.
  • Pressure testing on these channels when they are flowing.

hussam's update